BK Metals is currently operating at full capacity. The profit margin and the dividend payout ratio are held constant. Net working capital and fixed assets vary directly with sales. The company currently has current liabilities of $3,950, long-term debt of $14,700, net working capital of $7,850, net fixed assets of $27,600, owners' equity of $20,750, net income of $2,900, and dividends paid of $870. What is the external financing need if sales increase by 11 percent?

A -768

B. -145

C.972

D. 896

E. 1646
Answer #1

As current assets = current liabilities + net working capital, also,

Net working capital and fixed assets vary directly with sales,

If sales increase by 11 percent then, current assets will be

=110%(Net working capital) + current liabilities

=110%(7850) + 3950

=12663.5

If sales increase by 11 percent then, fixed assets will be

=110%(fixed assets)

=110%(27600)

=30636

Net income = 110%(2900) = 3219

New dividend = 110%(870) = 965.7

Net transfer to balance sheet = Net income - dividend = 3219 - 965.7 = 2253.3

Total assets = current assets + fixed assets = 12663.5 + 30636 = 43299.5

Shareholders fund = Owners equity + Net transer = 20750 + 2253.3 = 23003.3

Liabilities = Current liabilities + Debt = 3950 + 14700 = 18650

As assets = shareholders fund + liabilities

43299.5 = 23003.3 + 18650 + external financing

external financing = 1646.2 or 1646

The answer is e.